# TranSimHub: A Unified Air–Ground Simulation Platform for Multi-Modal Perception and Decision-Making

### v1

@@ Abstract
Air–ground collaborative intelligence is becoming a key approach for next-generation urban intelligent transportation management, where aerial and ground systems work together on perception, communication, and decision-making. However, the lack of a unified multi-modal simulation environment has limited progress in studying cross-domain perception, coordination under communication constraints, and joint decision optimization. To address this gap, we present TranSimHub, a unified simulation platform for air–ground collaborative intelligence. TranSimHub offers synchronized multi-view rendering across RGB, depth, and semantic segmentation modalities, ensuring consistent perception between aerial and ground viewpoints. It also supports information exchange between the two domains and includes a causal scene editor that enables controllable scenario creation and counterfactual analysis under diverse conditions such as different weather, emergency events, and dynamic obstacles. We release TranSimHub as an open-source platform that supports end-to-end research on perception, fusion, and control across realistic air and ground traffic scenes. Our code is available at https://github.com/Traffic-Alpha/TranSimHub.
@@ 1 Introduction
Although significant progress has been made in both ground-based and aerial simulation research, existing platforms remain isolated and domain-specific. Ground simulators such as VISSIM [1], Aimsun [2], and SUMO [3] have advanced studies in autonomous driving and traffic signal control, yet they focus mainly on road-level interactions among vehicles, traffic lights, and pedestrians, without incorporating aerial perception or communication. Conversely, aerial simulators such as AirSim [4] $Fe^{3}$ [5] and PX4 [6] specialize in flight control, path planning, and visual perception for UAVs or UAMs but lack modeling of ground dynamics or infrastructure coordination. Consequently, researchers currently lack a unified framework to investigate multi-agent cooperation, cross-domain perception fusion, and decision-making under communication constraints—capabilities essential for large-scale, intelligent air–ground mobility systems.
@@ 2 System Overview of TranSimHub
The Integration Interface Layer bridges TranSimHub with external ecosystems through standardized and extensible APIs. A Gym-compatible interface ensures seamless interoperability with reinforcement learning frameworks such as Stable Baselines3 [7] and TorchRL [8], supporting both single-agent and multi-agent learning paradigms. (M)LLM-driven control is enabled through LangChain [9], facilitating natural language–based interaction with simulated agents. For high-fidelity visualization, TranSimHub leverages Blender [10] for more realistic rendering and also supports Panda3D [11] for high efficiency simulation. In addition, GNS3 [12] and WinProp [13] are integrated to model inter-entity communication and wireless propagation, enabling comprehensive studies on perception, coordination, and decision-making across air–ground collaborative systems.
@@ 3.1 Ground Simulators
Ground-based traffic simulators have long served as the cornerstone of urban mobility research, offering quantitative tools to model vehicle interactions, multi-modal flows, and network-level dynamics. Early traffic assignment systems such as DynaMIT [14] and DynaSmart [15] enabled the analysis of dynamic routing and congestion propagation under varying demand conditions, laying the conceptual groundwork for subsequent microscopic and mesoscopic simulation frameworks. However, these early systems were often constrained by limited flexibility in behavioral modeling and scalability across city-scale networks.
Building upon these foundations, modern simulators such as VISSIM [1], Aimsun [2], and SUMO [3] have become the standards for high-fidelity ground traffic modeling. VISSIM and Aimsun emphasize detailed car-following and lane-changing mechanisms while supporting corridor- to region-scale simulations with customizable signal control logic. In contrast, SUMO, as a fully open-source platform, prioritizes extensibility and reproducibility, facilitating integration with reinforcement learning and optimization pipelines for adaptive traffic management. Together, these simulators provide versatile foundations for analyzing congestion, traffic signal coordination, and multi-modal interactions within complex urban environments.
As simulation requirements have expanded from flow-level dynamics to human-centric behavior, the focus has increasingly shifted toward agent-based and activity-based frameworks such as MATSim [16], BEAM [17], and POLARIS [18], which simulate individual decision-making, mode choice, and temporal scheduling at large scales. These models bridge microscopic realism with macroscopic scalability, enabling system-level analyses of transport policy, energy consumption, and social equity.
Meanwhile, specialized simulators have emerged to support distinct research domains. CARLA [19], MetaDrive [20], SMARTS [21] and Flow [22] provide photorealistic and physics-based environments for autonomous driving and control research, while CityFlow [23], CityFlowER [24] and Libsignal [25] focus specifically on large-scale traffic signal optimization with efficient parallel architectures.
@@ 3.2 Air Simulators
Airspace simulation platforms have evolved in parallel with ground-based systems, forming the analytical foundation for research on air traffic management and UAM. Early simulators such as FACET [26] and ATM TestBed [27], developed by NASA, focused on trajectory-level modeling of aircraft operations within the National Airspace System (NAS), enabling large-scale evaluations of traffic flow management, sector capacity, and conflict detection. However, these systems were primarily designed for conventional aviation and thus lack the scalability and flexibility required to simulate dense, low-altitude UAM environments. To address these limitations, a new generation of airspace simulators has been developed with an emphasis on flexibility, scalability, and modularity. Fe³ [5] provides GPU-accelerated simulations for high-density low-altitude flight scenarios, supporting rapid evaluation of separation assurance and contingency management strategies. VertiSim [28] models vertiport-level operations such as layout design, pad scheduling, and resource allocation, offering a practical environment for assessing takeoff and landing throughput constraints. Meanwhile, DTALite [29] and the open-source BlueSky [30] extend air traffic simulation to metropolitan and national scales, enabling trajectory optimization and concept-of-operations validation for emerging UAM networks. Collectively, these simulators represent a shift from traditional air traffic modeling toward integrated, multi-scale frameworks that can accommodate future UAM operations.
Beyond operational modeling, several simulators have been designed to support environmental, policy, and perception-oriented research, further broadening the scope of airspace simulation. AEDT [31] and AEIC [32] integrate noise, emission, and performance models to quantify the ecological footprint of aviation activities, while AirTraf 2.0 [33] couples flight trajectory simulation with the EMAC atmospheric chemistry model to study contrail formation and climate effects. In parallel, new simulators such as OpenUAV [34] and OpenFly [35] extend airspace simulation into the vision-language navigation domain, allowing aerial agents to perform perception-grounded reasoning and goal-directed navigation in photorealistic 3D environments. Together, these platforms form a layered and interdisciplinary simulation ecosystem—from vertiport micro-operations to macro-scale environmental analyses—laying the groundwork for integrating multimodal perception, language understanding, and decision-making into next-generation UAM and air–ground collaborative systems.

### v2

@@ Abstract
Air–ground collaborative intelligence is becoming a key approach for next-generation urban intelligent transportation management, where aerial and ground systems work together on perception, communication, and decision-making. However, the lack of a unified multi-modal simulation environment has limited progress in studying cross-domain perception, coordination under communication constraints, and joint decision optimization. To address this gap, we present TranSimHub, a unified simulation platform for air–ground collaborative intelligence. TranSimHub offers synchronized multi-view rendering across RGB, depth, and semantic segmentation modalities, ensuring consistent perception between aerial and ground viewpoints. It also supports information exchange between the two domains and includes a causal scene editor that enables controllable scenario creation and counterfactual analysis under diverse conditions such as different weather, emergency events, and dynamic obstacles. We release TranSimHub as an open-source platform that supports end-to-end research on perception, fusion, and control across realistic air and ground traffic scenes. Our code is available at https://github.com/Traffic-Alpha/TransSimHub.
@@ 1 Introduction
Although significant progress has been made in both ground-based and aerial simulation research, existing platforms remain isolated and domain-specific. Ground simulators such as VISSIM [fellendorf2010microscopic], Aimsun [casas2010traffic], and SUMO [SUMO2018] have advanced studies in autonomous driving and traffic signal control, yet they focus mainly on road-level interactions among vehicles, traffic lights, and pedestrians, without incorporating aerial perception or communication. Conversely, aerial simulators such as AirSim [shah2017airsim] $Fe^{3}$ [xue2018fe3] and PX4 [meier2015px4] specialize in flight control, path planning, and visual perception for UAVs or UAMs but lack modeling of ground dynamics or infrastructure coordination. Consequently, researchers currently lack a unified framework to investigate multi-agent cooperation, cross-domain perception fusion, and decision-making under communication constraints—capabilities essential for large-scale, intelligent air–ground mobility systems.
@@ 2 System Overview of TranSimHub
The Integration Interface Layer bridges TranSimHub with external ecosystems through standardized and extensible APIs. A Gym-compatible interface ensures seamless interoperability with reinforcement learning frameworks such as Stable Baselines3 [raffin2021stable] and TorchRL [bou2023torchrl], supporting both single-agent and multi-agent learning paradigms. (M)LLM-driven control is enabled through LangChain [Chase_LangChain_2022], facilitating natural language–based interaction with simulated agents. For high-fidelity visualization, TranSimHub leverages Blender [Blender2022] for more realistic rendering and also supports Panda3D [goslin2004panda3d] for high efficiency simulation. In addition, GNS3 [neumann2015book] and WinProp [hoppe2017wave] are integrated to model inter-entity communication and wireless propagation, enabling comprehensive studies on perception, coordination, and decision-making across air–ground collaborative systems.
@@ 3.1 Ground Simulators
Ground-based traffic simulators have long served as the cornerstone of urban mobility research, offering quantitative tools to model vehicle interactions, multi-modal flows, and network-level dynamics. Early traffic assignment systems such as DynaMIT [ben2002real] and DynaSmart [mahmassani2001dynamic] enabled the analysis of dynamic routing and congestion propagation under varying demand conditions, laying the conceptual groundwork for subsequent microscopic and mesoscopic simulation frameworks. However, these early systems were often constrained by limited flexibility in behavioral modeling and scalability across city-scale networks.
Building upon these foundations, modern simulators such as VISSIM [fellendorf2010microscopic], Aimsun [casas2010traffic], and SUMO [SUMO2018] have become the standards for high-fidelity ground traffic modeling. VISSIM and Aimsun emphasize detailed car-following and lane-changing mechanisms while supporting corridor- to region-scale simulations with customizable signal control logic. In contrast, SUMO, as a fully open-source platform, prioritizes extensibility and reproducibility, facilitating integration with reinforcement learning and optimization pipelines for adaptive traffic management. Together, these simulators provide versatile foundations for analyzing congestion, traffic signal coordination, and multi-modal interactions within complex urban environments.
As simulation requirements have expanded from flow-level dynamics to human-centric behavior, the focus has increasingly shifted toward agent-based and activity-based frameworks such as MATSim [w2016multi], BEAM [gopal2017modeling], and POLARIS [auld2016polaris], which simulate individual decision-making, mode choice, and temporal scheduling at large scales. These models bridge microscopic realism with macroscopic scalability, enabling system-level analyses of transport policy, energy consumption, and social equity.
Meanwhile, specialized simulators have emerged to support distinct research domains. CARLA [dosovitskiy2017carla], MetaDrive [li2022metadrive], SMARTS [pmlr-v155-zhou21a] and Flow [wu2021flow] provide photorealistic and physics-based environments for autonomous driving and control research, while CityFlow [zhang2019cityflow], CityFlowER [da2024cityflower] and Libsignal [mei2024libsignal] focus specifically on large-scale traffic signal optimization with efficient parallel architectures.
@@ 3.2 Air Simulators
Airspace simulation platforms have evolved in parallel with ground-based systems, forming the analytical foundation for research on air traffic management and UAM. Early simulators such as FACET [bilimoria2001facet] and ATM TestBed [robinson2017development], developed by NASA, focused on trajectory-level modeling of aircraft operations within the National Airspace System (NAS), enabling large-scale evaluations of traffic flow management, sector capacity, and conflict detection. However, these systems were primarily designed for conventional aviation and thus lack the scalability and flexibility required to simulate dense, low-altitude UAM environments. To address these limitations, a new generation of airspace simulators has been developed with an emphasis on flexibility, scalability, and modularity. Fe³ [xue2018fe3] provides GPU-accelerated simulations for high-density low-altitude flight scenarios, supporting rapid evaluation of separation assurance and contingency management strategies. VertiSim [yedavalli2021simuam] models vertiport-level operations such as layout design, pad scheduling, and resource allocation, offering a practical environment for assessing takeoff and landing throughput constraints. Meanwhile, DTALite [zhou2014dtalite] and the open-source BlueSky [hoekstra2016bluesky] extend air traffic simulation to metropolitan and national scales, enabling trajectory optimization and concept-of-operations validation for emerging UAM networks. Collectively, these simulators represent a shift from traditional air traffic modeling toward integrated, multi-scale frameworks that can accommodate future UAM operations.
Beyond operational modeling, several simulators have been designed to support environmental, policy, and perception-oriented research, further broadening the scope of airspace simulation. AEDT [roof2007aviation] and AEIC [simone2013rapid] integrate noise, emission, and performance models to quantify the ecological footprint of aviation activities, while AirTraf 2.0 [yamashita2020newly] couples flight trajectory simulation with the EMAC atmospheric chemistry model to study contrail formation and climate effects. In parallel, new simulators such as OpenUAV [gao2025openfly] and OpenFly [wang2024towards] extend airspace simulation into the vision-language navigation domain, allowing aerial agents to perform perception-grounded reasoning and goal-directed navigation in photorealistic 3D environments. Together, these platforms form a layered and interdisciplinary simulation ecosystem—from vertiport micro-operations to macro-scale environmental analyses—laying the groundwork for integrating multimodal perception, language understanding, and decision-making into next-generation UAM and air–ground collaborative systems.